# NoisyNN: Exploring the Impact of Information Entropy Change in Learning Systems

### 18 Dec 2023 v2

@@ 1 $\times$ 1 average pool, 1000-d fc, softmax
Appendix C Experimental Setting
We introduce the implementation details in this part. Model details are shown in Table 8 and 9. The image resolution is $224\times 224$ for all the experiments. Pre-trained models on ImageNet are used as the backbone. We train all ResNet and ViT-based models using AdamW optimizer [34]. We set the learning rate of each parameter group using a cosine annealing schedule with a minimum of $71e-$ 7. The data augmentation for training only includes the random resized crop and normalization.
@@ Appendix C Experimental Setting
Appendix D More Experiment Results
@@
D.1 Inject Positive Noise to Variants of ViT
@@
D.2 Positive Noise on TinyImageNet
@@
D.3 Positive Noise for Domain Adaptation
@@
Recently, transformer-based methods achieved SOTA results on UDA, therefore, we evaluate the ViT-B with the positive noise on widely used UDA benchmarks. Here the positive noise is the linear transform noise identical to that used in the classification task. The positive noise is injected into the last layer of the model, the same as the classification task. The datasets include Office Home [61] and VisDA2017 [42]. Office-Home[61] has 15,500 images of 65 classes from four domains: Artistic (Ar), Clip Art (Cl), Product (Pr), and Real-world (Rw) images. VisDA2017 is a Synthetic-to-Real object recognition dataset, with more than 0.2 million images in 12 classes. We use the ViT-B with a $16\times 16$ patch size, pre-trained on ImageNet. We use minibatch Stochastic Gradient Descent (SGD) optimizer [45] with a momentum of 0.9 as the optimizer. The batch size is set to 32. We initialized the learning rate as 0 and linearly warm up to 0.05 after 500 training steps. The results are shown in Table 13 and 14. The methods above the black line are based on CNN architecture, while those under the black line are developed from the Transformer architecture. The NoisyTVT-B, i.e., TVT-B with positive noise, achieves better performance than existing works. These results show that positive noise can improve model generality, therefore, benefit deep models in domain adaptation tasks.

### 02 Feb 2024 v3

@@ C.1.1 Optimal Quality Matrix of Linear Transform Noise
C.2 Add Linear Transform Noise to Raw Images
In this case, the task entropy with linear transform noise can be formulated as:
where $I$ is an identity matrix, and $Q$ is derived from $I$ using elementary row operations. Assuming that the pixels of the raw images follow a Gaussian distribution. The conditional distribution of $\boldsymbol{Y}$ given $\boldsymbol{X}+Q\boldsymbol{X}$ is also multivariate subjected to the normal distribution, which can be formulated as:
Since the linear transform matrix is invertible, applying the linear transform to $\boldsymbol{X}$ does not alter the distribution of the $\boldsymbol{X}$. It is straightforward to obtain:
The entropy change of 0 indicates that the implementation of linear transformation to the raw images could not help reduce the complexity of the task.
Appendix D Influence of Salt-and-pepper Noise on Task Entropy
D.1 Inject Salt-and-pepper Noise in Latent Space
$\begin{split}\bigtriangleup S(\mathcal{T},Q\boldsymbol{Z})=&H(\mathcal{T};\boldsymbol{Z})-H(\mathcal{T};\boldsymbol{Z}\boldsymbol{\epsilon})\\ =&H(\boldsymbol{Y};\boldsymbol{Z})-H(\boldsymbol{Z})-(H(\boldsymbol{Y};\boldsymbol{Z}\boldsymbol{\epsilon})-H(\boldsymbol{Z}))\\ =&H(\boldsymbol{Y};\boldsymbol{Z})-H(\boldsymbol{Y};\boldsymbol{Z}\boldsymbol{\epsilon})\\ =&-\sum_{\boldsymbol{Z}\in\mathcal{Z}}\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{Z},\boldsymbol{Y})\log p(\boldsymbol{Z},\boldsymbol{Y})+\sum_{\boldsymbol{Z}\in\mathcal{Z}}\sum_{\boldsymbol{Y}\in\mathcal{Y}}\sum_{\boldsymbol{\epsilon}\in\mathcal{E}}p(\boldsymbol{Z\epsilon},\boldsymbol{Y})\log p(\boldsymbol{Z\epsilon},\boldsymbol{Y})\\ =&\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{Z},\boldsymbol{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{Z\epsilon},\boldsymbol{Y})}\right]\\ =&\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{Z},\boldsymbol{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{Z},\boldsymbol{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{\epsilon})}\right]\\ =&-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{\epsilon})}\right]\\ =&-H(\boldsymbol{\epsilon})\end{split}$ (89)
The entropy change is smaller than 0, therefore, the salt-and-pepper is a pure detrimental noise to the learning task.
D.2 Add Salt-and-pepper Noise to Raw Images
The task entropy with salt-and-pepper noise is rewritten as:
where
Therefore, the entropy change with salt-and-pepper noise is:
$\bigtriangleup S(\mathcal{T},Q\boldsymbol{X})=H(\mathcal{T};\boldsymbol{X})-H(\mathcal{T};\boldsymbol{X}\boldsymbol{\epsilon})=0$ (93)
Figure 3: Differnet Noise Injection Modes. The green tick mark means noise injection operations are implemented, while the red cross means no noise injection. We adopt the Mode 1 in this work. Modes 2 and 3 are for future exploration.
Appendix E Experimental Setting
We introduce the implementation details in this part. The noise was added to both the training stage and the inference stage. Model details are shown in Table 8 and 9. The image resolution is $224\times 224$ for all the experiments. Pre-trained models on ImageNet-21K are used as the backbone. We train all ResNet and ViT-based models using AdamW optimizer [35]. We set the learning rate of each parameter group using a cosine annealing schedule with a minimum of $71e-$ 7. The data augmentation for training only includes the random resized crop and normalization.
@@ 1 $\times$ 1 average pool, 1000-d fc, softmax
Appendix F More Experiment Results
@@
F.1 Inject Positive Noise to Variants of ViT
@@
F.2 Positive Noise on TinyImageNet
@@
F.3 Positive Noise for Domain Adaptation
@@
Recently, transformer-based methods achieved SOTA results on UDA, therefore, we evaluate the ViT-B with the positive noise on widely used UDA benchmarks. Here the positive noise is the linear transform noise identical to that used in the classification task. The positive noise is injected into the last layer of the model, the same as the classification task. The datasets include Office Home [62] and VisDA2017 [43]. Office-Home[62] has 15,500 images of 65 classes from four domains: Artistic (Ar), Clip Art (Cl), Product (Pr), and Real-world (Rw) images. VisDA2017 is a Synthetic-to-Real object recognition dataset, with more than 0.2 million images in 12 classes. We use the ViT-B with a $16\times 16$ patch size, pre-trained on ImageNet. We use minibatch Stochastic Gradient Descent (SGD) optimizer [46] with a momentum of 0.9 as the optimizer. The batch size is set to 32. We initialized the learning rate as 0 and linearly warm up to 0.05 after 500 training steps. The results are shown in Table 13 and 14. The methods above the black line are based on CNN architecture, while those under the black line are developed from the Transformer architecture. The NoisyTVT-B, i.e., TVT-B with positive noise, achieves better performance than existing works. These results show that positive noise can improve model generality and, therefore, benefit deep models in domain adaptation tasks.